Shown below is a ray of light going from air through crown glass into water, such as going into a fish tank.

1 2 02 3

Calculate the amount the ray is displaced by the glass (Δx), given that the incident angle is θ1=34 deg and the glass is 1.1 cm thick.
Answer #1

Solution)

N2/n1= sin theta1/sin theta 2

Sin theta 2= sin 34/1.52

Theta2= 21.58 degrees

Tan theta1=x1/1.1

X1= 0.74 cm

Than theta2=x2/1.1

X2= 0.435 cm

Now, delta x= x1-x2 =0.74-0.435 =0.304 cm (Ans)

======

Good luck!:)